OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

A. y = −4x + 20 B. y = 4x + 20 C. y = −4x – 35 D. y = 4x – 35 E. y = -4x + 4

OpenStudy (anonymous):

the standard form is: y−y1=m(x−x1) where m is your slope and y_1 and x_1 are your points

OpenStudy (jhannybean):

\[\large m=-4 \ , \ (-8,-3)\] Use slope intercept form \[\large y=mx+b\]\[\large -3=(-4)(-8)+b\] Solve for your y-intercept first - "b" \[\large -3 = 32 +b\]\[\large -3-32=b\]\[\large b = -35\]Now that we have our slope and y intercept, we can write our equation with our new found values into slope intercept form.\[\large y= -4x-35\]